OpenStudy (anonymous):

Solve for x. 8 over x equals x over 11

OpenStudy (anonymous):

8 over x equals x over 11. 8/x = x/11 8(11)=x^2 x=sqrt(88) x=2sqrt22 ans or 2 radical22

OpenStudy (anonymous):

\[\frac{ 8 }{ x }=\frac{ x }{ 11 }\] \[x^2=88\] Can you solve it from there?

OpenStudy (anonymous):

You're not supposed to give them the answer...

OpenStudy (anonymous):

thank u xoya and jay567

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

Yw :)